• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name X. Zheng/xfzhu

    Log entry time 20:46:57 on July30,2001

    Entry number 68389

    keyword=Target - Polarization measurement and configuration change

    19:56 stop data taking at (in, out), move target to pick-up
    19:57 NMR with 3 lasers, P=41.42%
    19:59 EPR with 3 lasers, P=42.4%
    20:05 turn off 3 longitudinal lasers,
    20:06 Rotate field to 180, move target half wave plates in
    20:15 NMR with all lsers off, P=40.57%
    20:17~18, turn on 3 longitudinal lasers
    20:19 move target to pol 3he cell position
    20:27 EPR with 3 lasers on, P= 40.2%


    19:59 EPR configuration:
    holding field 0deg (L=-7.0573 T=-0.0870A)
    3 long lasers on (5,6,7), target at pickup coil position
    0th order fit:
    freq1: -46.58 kHz +/- 1.31 kHz | Polarisation: -42.54 % +/- 1.19 %
    freq2: 45.85 kHz +/- 1.31 kHz | Polarisation: 41.87 % +/- 1.2 %
    1st order fit
    freq1: -46.84 kHz +/- 1.3 kHz | Polarisation: -42.78 % +/- 1.19 %
    freq2: 46.45 kHz +/- 1.18 kHz | Polarisation: 42.42 % +/- 1.08 %

    20:27 EPR configuration:
    holding field 180, 3 long. lasers on, target at pol 3he cell position
    0th order fitting:
    freq1: -43.61 kHz +/- 1.39 kHz | Polarisation: -39.83 % +/- 1.27 %
    freq2: 44.11 kHz +/- 1.34 kHz | Polarisation: 40.28 % +/- 1.22 %

    1st order fitting:
    freq1: -44.02 kHz +/- 1.27 kHz | Polarisation: -40.2 % +/- 1.16 %
    freq2: 44.28 kHz +/- 1.33 kHz | Polarisation: 40.44 % +/- 1.22 %